OpenStudy (anonymous):

At an archery competiton, Jamie and Marshall's team tied Shawna and Wyatt's team for the top scores. Marshall scored 12 less than Jamie. Shawna scored twice as much as Marshall. Wyatt scored 4 less than Shawna. What was each person's individual score?

OpenStudy (anonymous):

Let Jamie's score be J, Marshall's be M and Shawna's be S and Wyatt's be W

OpenStudy (anonymous):

So now from the question M is 12 less than J, so can we write the following equation? M=J-12?

OpenStudy (anonymous):

m=j-12

OpenStudy (anonymous):

And can we write S=2M?

OpenStudy (anonymous):

and W=S-4?

OpenStudy (anonymous):

s=2m w=s-4

OpenStudy (anonymous):

and finally as the combined scores of J and M are tied with those of S and W then, J+M = S+W right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

As m = j-12 so we can write it as m+12 = j ??

OpenStudy (anonymous):

As J+M = W+S so (M+12)+M = S+W but S=2M and W = S-4 = 2M -4 so now (M+12)+M = 2M + 2M-4

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So what is the value of M?

OpenStudy (anonymous):

is it 8?

OpenStudy (anonymous):

yup exactly!

OpenStudy (anonymous):

and J,W and S?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

j=20 w=12 s=16 m=8 correct?

OpenStudy (anonymous):

absolutely!

OpenStudy (anonymous):

Thank you very much